A Seller sells his product on 14% profit of the market price. If the selling price is 3420 then find the
market price. 
please give me formula  with properly solved with calculations

Grade:12

1 Answers

Arun
4 years ago
 
Let the Market price be X.
A gain is seen in the given problem, hence,
Selling price = Market price + Profit
3420= X +14% of 3420
X= 3420 - 14% of 3420
X= 3420 - 478.8
 
Hope it helps
 
Regards
Arun (askIITians forum expert)
X= 2941.2